OpenStudy (nannie):

Would (a+7)² or a²+49 be the best answer to a²+14a+49, for completly factoring?

OpenStudy (anonymous):

(a+7)^2 = a^2 + 2*7*a + 7^2 = a^2 + 14a + 49

OpenStudy (nannie):

TY again!!!!

OpenStudy (anonymous):

coolsector has the correct answer, however turning \[(a+7)^2\] into \[a^2+14a+49\] is called "multiplying" not "factoring" "factoring" means rewriting \[a^2+14a+49\] as \[(a+7)^2\]